# Scalable Graph Generative Modeling via Substructure Sequences

### v1

@@ Abstract
Graph neural networks (GNNs) has been predominantly driven by message-passing, where node representations are iteratively updated via local neighborhood aggregation. Despite their success, message-passing suffers from fundamental limitations—including constrained expressiveness, over-smoothing, over-squashing, and limited capacity to model long-range dependencies. These issues hinder scalability: increasing data size or model size often fails to yield improved performance, limiting the viability of GNNs as backbones for graph foundation models. In this work, we explore pathways beyond message-passing and introduce Generative Graph Pattern Machine (G2PM), a generative Transformer pre-training framework for graphs. G2PM represents graph instances (nodes, edges, or entire graphs) as sequences of substructures, and employs generative pre-training over the sequences to learn generalizable, transferable representations. Empirically, G2PM demonstrates strong scalability: on the ogbn-arxiv benchmark, it continues to improve with model sizes up to 60M parameters, outperforming prior generative approaches that plateau at significantly smaller scales (e.g., 3M). In addition, we systematically analyze the model design space, highlighting key architectural choices that contribute to its scalability and generalization. Across diverse tasks—including node classification, graph classification, and transfer learning—G2PM consistently outperforms strong baselines, establishing a compelling foundation for scalable graph learning. The code and dataset are available at https://github.com/Zehong-Wang/G2PM.
@@ 1 Introduction
Transformer architectures [54] have emerged as the cornerstone of modern foundation models [7, 12, 2, 52]. Enabled by generative pre-training on massive unlabeled corpora [7, 11], Transformers learn transferable representations that can be efficiently adapted to diverse downstream tasks. This paradigm shift has redefined the landscape of machine learning, powering state-of-the-art systems in natural language processing and computer vision. Prominent examples include large language models (LLMs) [2, 52] and Vision Transformers (ViTs) [12], which showcase the scalability and generalization strength of this approach.
@@
Despite the transformative success of generative Transformer pre-training in text and vision domains, this paradigm has yet to bring comparable breakthroughs in the graph domain, posing a significant barrier to the development of graph foundation models [35, 60, 62]. Most existing approaches to graph pre-training remain grounded in message-passing graph neural networks (MPNNs) [28, 55, 15, 17, 34], which are known to suffer from fundamental limitations: constrained expressive power [72], over-smoothing [44], over-squashing [51], and poor capacity for modeling long-range dependencies [42]. These challenges significantly limit the scalability of MPNNs [38], where increasing the size of the model or training data does not reliably lead to improved performance. As a result, the emergence of scaling laws in graph learning—a critical property in the success of foundation models [25]—remains elusive. Moreover, current graph pre-training techniques are predominantly based on contrastive learning [41, 78, 59, 58, 48, 79, 70, 71, 40, 33], which has been shown to be less capable of learning generalizable semantic representations compared to generative objectives [19, 11, 49, 5, 7]. This reliance further compounds the scalability bottleneck of graph neural networks (GNNs).
@@
Figure 2: Model scaling behavior on the ogbn-arxiv with linear probing. G2PM consistently improves as model parameters increase, achieving 72.31% accuracy with 60M parameters—surpassing the current SOTA generative graph pre-training method [22]. In contrast, GraphMAE [21] and BGRL [48] exhibit performance degradation, indicating limited scalability.
@@
Figure 3: Data scaling behavior on the ogbn-arxiv and ogbn-products with linear probing. G2PM demonstrates robust improvements with more pre-training data, reflecting superior scalability. In contrast, GraphMAE [21] and BGRL [48] peak at small data ratios and degrade with more data, suggesting overfitting, poor regularization, or inefficient data utilization.
@@
In this work, we aim to extend the success of Transformer-based generative pre-training to the graph domain, with the goal of enabling scalable graph representation learning. We consider the effectiveness of Transformers stems from a common principle: Transformer-based models represent modality-specific instances using sequences of high-level semantic tokens, and apply generative objectives for pre-training. To realize this paradigm for graphs, we begin by identifying three fundamental challenges that distinguish graph-structured data from Euclidean modalities such as text and images. (1) Absence of Sequence Structure: Unlike text or images, which naturally possess ordered or grid-like sequences compatible with Transformer training, graphs lack a sequence (no matter ordered or unordered) for nodes, edges, or subgraphs. This complicates the adaptation of Transformer. (2) Semantic Granularity: Graph elements—such as nodes or edges—typically encode low-level semantics, while tokens in language (i.e., words) or vision (i.e., patches) often correspond to higher-level concepts. It is unclear whether Transformers can effectively learn from such low-level representations in a generative fashion. (3) Scalability Bottlenecks: Existing Graph Transformers (GTs) [29, 9, 13] often treat individual nodes as tokens and focus on pairwise relationships, leading to sequence lengths that scale with the number of nodes. Due to the quadratic time complexity of Transformers [26], this design choice limits scalability to large graphs, confining GTs primarily to small graphs such as molecular graphs [42].
@@
  • We validate G2PM across multiple benchmarks, showing state-of-the-art performance on node and graph classification, along with strong generalization in cross-graph transfer tasks.
@@
2 Method: Generative Graph Pattern Machine
@@ 2.1 Graph Representations
Our core idea is to represent any graph instance (node, link, or graph) as a sequence of substructures. In natural language, sequences are constructed using a predefined vocabulary of words or subwords; in vision, raster-scan tokenizers divide images into patch sequences. However, extending this notion of tokenization to graphs poses two key challenges. First, it is non-trivial to define a universal vocabulary of graph substructures, as the semantic relevance of patterns is highly domain-specific. For example, triangle motifs are prominent in social networks, whereas ring structures are more meaningful in molecular graphs. Second, tokenizing graphs by substructure matching incurs high computational cost: subgraph isomorphism is NP-complete [14], rendering it impractical for large-scale graphs—especially when compared to the linear-time tokenization of text and images [46, 12].
Tokenizer. To overcome these challenges, we adopt a random walk-based tokenizer that samples substructure patterns on the fly, bypassing the need for a predefined vocabulary. This strategy offers an efficient and scalable tokenization mechanism [63, 76]. As illustrated in Figure 4(b), a random walk such as $[1,2,3,1,4,3]$ corresponds to a diamond-shaped substructure. Formally, we define an unbiased random walk $w$ of fixed length $L$ as a sequence sampled from a Markov chain:
@@
For each graph instance, we sample $k$ random walks to serve as substructure patterns. To balance locality and global context, we leverage unbiased sampling strategies [39, 63]. Rather than constructing explicit subgraphs from these walks, we treat each as a node sequence and directly encode it using a Transformer, which has been shown to effectively capture structural inductive biases in graphs [63].
@@
Figure 4: Overview of G2PM pre-training. (a) Given a graph instance, we first apply a random walk-based tokenizer to extract substructure patterns. Some patterns are randomly masked and the visible patterns are fed into a G2PM encoder. The encoder outputs are concatenated with special mask tokens and passed to a G2PM decoder to reconstruct the masked substructures. (b) We tokenize substructures by performing random walks over the graph, where each walk represents a substructure, which is proved to be effective [63]. The substructure embeddings are modeled via Transformer.
Eliminating Message-Passing Bottlenecks. Our tokenizer is entirely message-passing-free, allowing G2PM to bypass key limitations of traditional GNNs. Notably, the random walk-based representations in G2PM effectively capture long-range dependencies [63, 24], exceed the expressiveness of 1-WL tests [50, 64, 37], and mitigate over-smoothing and over-squashing effects [56].
No Positional Embeddings. Unlike standard Transformers, which depend on positional encodings, we find that adding position information to node sequences offers no consistent benefit (Table 1) and incurs cubic time complexity [29]. We thus omit positional embeddings entirely, simplifying the architecture without compromising performance.
@@ 2.2 G2PM Backbone
We adopt the standard Transformer architecture [54] as the backbone for both G2PM encoder and decoder. The input to the Transformer is a sequence of graph substructure tokens ${\mathbf{P}}=[{\mathbf{p}}_{1},{\mathbf{p}}_{2},\dots,{\mathbf{p}}_{n}]$, where each ${\mathbf{p}}_{i}$ denotes the embedding of a sampled substructure. A single Transformer layer operates via a combination of self-attention and feed-forward networks, formally defined as:
@@
where ${\mathbf{W}}_{{\mathbf{Q}}},{\mathbf{W}}_{{\mathbf{K}}},{\mathbf{W}}_{{\mathbf{V}}}$ are learnable projection matrices and $d_{\text{out}}$ denotes the dimensionality of the output embeddings. The FFN is implemented as a two-layer multilayer perceptron with non-linearity. Following standard practice, we employ multi-head self-attention [54], where multiple attention mechanisms are applied in parallel and their outputs concatenated. We stack multiple Transformer layers to enhance model capacity. We denote the output of the final ($L$-th) Transformer layer as ${\mathbf{P}}^{L}=[{\mathbf{p}}^{L}_{1},{\mathbf{p}}^{L}_{2},\dots,{\mathbf{p}}^{L}_{n}]$, representing the learned contextualized embeddings of graph substructures.
@@ 2.3 Pre-Training G2PM: Masked Substructure Modeling
To pre-train the G2PM model, we introduce a generative Transformer objective masked substructure modeling (MSM). Given a sequence of substructures extracted from a graph, we randomly mask a subset and train the model to reconstruct the masked substructures conditioned on the visible ones. This process follows the conditional factorization principle used in masked language modeling [11]:
@@
The key intuition is that meaningful graph substructures exhibit strong interdependencies—such as hierarchical relationships, functional reinforcement, or functional exclusion (Appendix A)—allowing certain substructures to be predictive of others. Concretely, given a tokenized substructure sequence ${\mathbf{P}}=[{\mathbf{p}}_{1},\dots,{\mathbf{p}}_{n}]$, we randomly retain $k$ visible tokens to form ${\mathbf{P}}_{\text{vis}}$, while masking the remaining $n-k$ tokens. The visible sequence ${\mathbf{P}}_{\text{vis}}$ is passed through the G2PM encoder to produce contextual embeddings ${\mathbf{H}}_{\text{vis}}$. We then insert learnable mask tokens at the masked positions to recover the full-length sequence, and feed the sequence into the G2PM decoder for reconstruction ${\mathbf{R}}=[{\mathbf{r}}_{1},\dots,{\mathbf{r}}_{n}]$:
@@
Augmenting Substructure Patterns. To improve robustness and representation quality, we adopt a mixed augmentation strategy inspired by corruption-agnostic pre-training [19, 66]. We define four augmentations in two categories: Feature-level—(1) feature masking (zeroing a subset of features) and (2) node masking (masking entire nodes); and Structure-level—(3) substructure corruption (dropping nodes within a walk) and (4) substructure injection (replacing walk nodes with random nodes). During training, we apply one feature-level and one structure-level augmentation per instance to promote generalization across diverse perturbations.
@@
(a) Model architecture. Our G2PM design (MAE-style [19]) outperforms SimMIM-style [66] pre-training.
@@ 3.1 Model Architecture
Model Dimension. We set the default hidden dimension to 768, with a 3-layer encoder and a 1-layer decoder. As shown in Table 1, increasing model size consistently improves performance, mirroring scaling trends observed in language and vision domains [7, 12]. Unlike message-passing models that often plateau with scale, G2PM continues to benefit from increased parameterization.
Model Layer. Table 5 shows that deeper encoders enhance performance, while increasing decoder capacity provides limited or negative returns, particularly under the MAE-style masked modeling framework. We attribute this to the simplicity of substructure reconstruction, where expressive decoders can lead to overfitting. Replacing the MAE-style design [19] with a SimMIM-style variant [66] further degrades performance. This suggests that MAE-style sparsity encourages the encoder to learn stronger context-aware representations, which better align with the inductive bias of graph data.
@@ 3.2 Reconstruction Target
Loss. We also evaluate the impact of loss functions. Replacing $\ell_{2}$ loss with $\ell_{1}$ leads to a notable drop in graph classification accuracy, suggesting that outlier-sensitive objectives better capture nuanced substructure semantics. Additionally, we incorporate anonymous walks [23] following Wang et al. [63], using a parallel head to reconstruct topological patterns. This auxiliary objective consistently boosts performance, reinforcing the utility of substructures as fundamental modeling units.

### v3

# Generative Graph Pattern Machine

@@ Abstract
Graph neural networks (GNNs) have been predominantly driven by message-passing, where node representations are iteratively updated via local neighborhood aggregation. Despite their success, message-passing suffers from fundamental limitations—including constrained expressiveness, over-smoothing, over-squashing, and limited capacity to model long-range dependencies. These issues hinder scalability: increasing data size or model size often fails to yield improved performance. To this end, we explore pathways beyond message-passing and introduce Generative Graph Pattern Machine (G2PM), a generative Transformer pre-training framework for graphs. G2PM represents graph instances (nodes, edges, or entire graphs) as sequences of substructures, and employs generative pre-training over the sequences to learn generalizable and transferable representations. Empirically, G2PM demonstrates strong scalability: on the ogbn-arxiv benchmark, it continues to improve with model sizes up to 60M parameters, outperforming prior generative approaches that plateau at significantly smaller scales (e.g., 3M). In addition, we systematically analyze the model design space, highlighting key architectural choices that contribute to its scalability and generalization. Across diverse tasks—including node/link/graph classification, transfer learning, and cross-graph pretraining—G2PM consistently outperforms strong baselines, establishing a compelling foundation for scalable graph learning. The code and dataset are available at https://github.com/Zehong-Wang/G2PM.
@@ 1 Introduction
Transformer architectures [vaswani2017attention] have emerged as the cornerstone of modern foundation models [brown2020language, dosovitskiy2021an, achiam2023gpt, touvron2023llama]. Enabled by generative pre-training on massive unlabeled corpora [brown2020language, devlin2019bert], Transformers learn transferable representations that can be efficiently adapted to diverse downstream tasks. This paradigm shift has redefined the landscape of machine learning, powering state-of-the-art systems in natural language processing and computer vision. Prominent examples include large language models (LLMs) [achiam2023gpt, touvron2023llama] and Vision Transformers (ViTs) [dosovitskiy2021an], which showcase the scalability and generalization strength of this approach.
@@
Despite the transformative success of generative Transformer pre-training in text and vision domains, this paradigm has yet to bring comparable breakthroughs in the graph domain, posing a significant barrier to the development of graph foundation models [mao2024graph, wang2024gft, wang2025git]. Most existing approaches to graph pre-training remain grounded in message-passing graph neural networks (MPNNs) [kipf2017semisupervised, velickovic2018graph, gilmer2017neural, hamilton2017inductive, ma2025adaptive], which are known to suffer from fundamental limitations: constrained expressive power [zhang2024beyond], over-smoothing [rusch2023survey], over-squashing [topping2021understanding], and poor capacity for modeling long-range dependencies [rampasek2022recipe]. These challenges significantly limit the scalability of MPNNs [morris2024position], where increasing the size of the model or training data does not reliably lead to improved performance. As a result, the emergence of scaling laws in graph learning—a critical property in the success of foundation models [kaplan2020scaling]—remains elusive. Moreover, current graph pre-training techniques are predominantly based on contrastive learning [qiu2020gcc, zhu2020deep, wang2024select, wang2023heterogeneous, thakoor2022largescale, zhu2021graph, you2020graph, you2021graph, qian2024dual, ma2023hypergraph], which has been shown to be less capable of learning generalizable semantic representations compared to generative objectives [he2022masked, devlin2019bert, tian2024visual, bao2021beit, brown2020language]. This reliance further compounds the scalability bottleneck of graph neural networks (GNNs).
@@
Figure 2: Model scaling behavior on the ogbn-arxiv with linear probing. G2PM consistently improves as model parameters increase, achieving 72.31% accuracy with 60M parameters—surpassing the current SOTA generative graph pre-training method [hou2023graphmae2]. In contrast, GraphMAE [hou2022graphmae] and BGRL [thakoor2022largescale] exhibit performance degradation, indicating limited scalability.
@@
Figure 3: Data scaling behavior on the ogbn-arxiv and ogbn-products with linear probing. G2PM demonstrates robust improvements with more pre-training data, reflecting superior scalability. In contrast, GraphMAE [hou2022graphmae] and BGRL [thakoor2022largescale] peak at small data ratios and degrade with more data, suggesting overfitting, poor regularization, or inefficient data utilization.
@@
In this work, we aim to extend the success of Transformer-based generative pre-training to the graph domain, with the goal of enabling scalable graph representation learning. We consider the effectiveness of Transformers stems from a common principle: Transformer-based models represent modality-specific instances using sequences of high-level semantic tokens, and apply generative objectives for pre-training. To realize this paradigm for graphs, we begin by identifying three fundamental challenges that distinguish graph-structured data from Euclidean modalities such as text and images. (1) Absence of Sequence Structure: Unlike text or images, which naturally possess ordered or grid-like sequences compatible with Transformer training, graphs lack a sequence (no matter ordered or unordered) for nodes, edges, or subgraphs. This complicates the adaptation of Transformer. (2) Semantic Granularity: Graph elements—such as nodes or edges—typically encode low-level semantics, while tokens in language (i.e., words) or vision (i.e., patches) often correspond to higher-level concepts. It is unclear whether Transformers can effectively learn from such low-level representations in a generative fashion. (3) Scalability Bottlenecks: Existing Graph Transformers (GTs) [kreuzer2021rethinking, chen2023nagphormer, dwivedi2020generalization] often treat individual nodes as tokens and focus on pairwise relationships, leading to sequence lengths that scale with the number of nodes. Due to the quadratic time complexity of Transformers [katharopoulos2020transformers], this design choice limits scalability to large graphs, confining GTs primarily to small graphs such as molecular graphs [rampasek2022recipe].
@@
  • We validate G2PM across multiple benchmarks, showing state-of-the-art performance on node/link/graph-level tasks, along with strong generalization in cross-graph transfer tasks.
@@
2 Generative Graph Pattern Machine
@@ 2.1 Graph Representations
Our core idea is to represent any graph instance (node, link, or graph) as a sequence of substructures. In natural language, sequences are constructed using a predefined vocabulary of words or subwords; in vision, raster-scan tokenizers divide images into patch sequences. However, extending this notion of tokenization to graphs poses two key challenges. First, it is non-trivial to define a universal vocabulary of graph substructures, as the semantic relevance of patterns is highly domain-specific. For example, triangle motifs are prominent in social networks, whereas ring structures are more meaningful in molecular graphs. Second, tokenizing graphs by substructure matching incurs high computational cost: subgraph isomorphism is NP-complete [garey2002computers], rendering it impractical for large-scale graphs—especially when compared to the linear-time tokenization of text and images [sennrich2015neural, dosovitskiy2021an].
Tokenizer. To overcome these challenges, we adopt a random walk-based tokenizer that samples substructure patterns on the fly, bypassing the need for a predefined vocabulary. This strategy offers an efficient and scalable tokenization mechanism [wang2025neural, zhao2022from]. As illustrated in Figure 4(b), a random walk such as $[1,2,3,1,4,3]$ corresponds to a diamond-shaped substructure. Formally, we define an unbiased random walk $w$ of fixed length $L$ as a sequence sampled from a Markov chain:
@@
For each graph instance, we sample $k$ random walks to serve as substructure patterns. To balance locality and global context, we leverage unbiased sampling strategies [perozzi2014deepwalk, wang2025neural]. Rather than constructing explicit subgraphs from these walks, we treat each as a node sequence and directly encode it using a Transformer, which has been shown to effectively capture structural inductive biases in graphs [wang2025neural].
@@
Figure 4: Overview of G2PM pre-training. (a) Given a graph instance, we first apply a random walk-based tokenizer to extract substructure patterns. Some patterns are randomly masked and the visible patterns are fed into a G2PM encoder. The encoder outputs are concatenated with special mask tokens and passed to a G2PM decoder to reconstruct the masked substructures. (b) We tokenize substructures by performing random walks over the graph, where each walk represents a substructure, which is proved to be effective [wang2025neural]. The substructure embeddings are modeled via Transformer.
Eliminating Message-Passing Bottlenecks. Our tokenizer is entirely message-passing-free, allowing G2PM to bypass key limitations of traditional GNNs. Notably, the random walk-based representations in G2PM effectively capture long-range dependencies [wang2025neural, jin2022raw], exceed the expressiveness of 1-WL tests [tonshoff2023walking, welke2023expectation, michel2023path], and mitigate over-smoothing and over-squashing effects [wang2024non].
No Positional Embeddings. Unlike standard Transformers, which depend on positional encodings, we find that adding position information to node sequences offers no consistent benefit (Table 1) and incurs cubic time complexity [kreuzer2021rethinking]. We thus omit positional embeddings entirely, simplifying the architecture without compromising performance.
@@ 2.2 G2PM Backbone
We adopt the standard Transformer architecture [vaswani2017attention] as the backbone for both G2PM encoder and decoder. The input to the Transformer is a sequence of graph substructure tokens ${\mathbf{P}}=[{\mathbf{p}}_{1},{\mathbf{p}}_{2},\dots,{\mathbf{p}}_{n}]$, where each ${\mathbf{p}}_{i}$ denotes the embedding of a sampled substructure. A single Transformer layer operates via a combination of self-attention and feed-forward networks, formally defined as:
@@
where ${\mathbf{W}}_{{\mathbf{Q}}},{\mathbf{W}}_{{\mathbf{K}}},{\mathbf{W}}_{{\mathbf{V}}}$ are learnable projection matrices and $d_{\text{out}}$ denotes the dimensionality of the output embeddings. The FFN is implemented as a two-layer multilayer perceptron with non-linearity. Following standard practice, we employ multi-head self-attention [vaswani2017attention], where multiple attention mechanisms are applied in parallel and their outputs concatenated. We stack multiple Transformer layers to enhance model capacity. We denote the output of the final ($L$-th) Transformer layer as ${\mathbf{P}}^{L}=[{\mathbf{p}}^{L}_{1},{\mathbf{p}}^{L}_{2},\dots,{\mathbf{p}}^{L}_{n}]$, representing the learned contextualized embeddings of graph substructures.
@@ 2.3 Pre-Training G2PM: Masked Substructure Modeling
To pre-train the G2PM model, we introduce a generative Transformer objective masked substructure modeling (MSM). Given a sequence of substructures extracted from a graph, we randomly mask a subset and train the model to reconstruct the masked substructures conditioned on the visible ones. This process follows the conditional factorization principle used in masked language modeling [devlin2019bert]:
@@
The key intuition is that meaningful graph substructures exhibit strong interdependencies—such as hierarchical relationships, functional reinforcement, or functional exclusion (Appendix A)—allowing certain substructures to be predictive of others. Concretely, given a tokenized substructure sequence ${\mathbf{P}}=[{\mathbf{p}}_{1},\dots,{\mathbf{p}}_{n}]$, we randomly retain $k$ visible tokens to form ${\mathbf{P}}_{\text{vis}}$, while masking the remaining nkn-k tokens. The visible sequence ${\mathbf{P}}_{\text{vis}}$ is passed through the G2PM encoder to produce contextual embeddings ${\mathbf{H}}_{\text{vis}}$. We then insert learnable mask tokens at the masked positions to recover the full-length sequence, and feed the sequence into the G2PM decoder for reconstruction ${\mathbf{R}}=[{\mathbf{r}}_{1},\dots,{\mathbf{r}}_{n}]$:
@@
Augmenting Substructure Patterns. To improve robustness and representation quality, we adopt a mixed augmentation strategy inspired by corruption-agnostic pre-training [he2022masked, xie2022simmim]. We define four augmentations in two categories: Feature-level—(1) feature masking (zeroing a subset of features) and (2) node masking (masking entire nodes); and Structure-level—(3) substructure corruption (dropping nodes within a walk) and (4) substructure injection (replacing walk nodes with random nodes). During training, we apply one feature-level and one structure-level augmentation per instance to promote generalization across diverse perturbations.
@@
(a) Model architecture. Our G2PM design (MAE-style [he2022masked]) outperforms SimMIM-style [xie2022simmim] pre-training.
@@ 3.1 Model Architecture
Model Dimension. We set the default hidden dimension to 768, with a 3-layer encoder and a 1-layer decoder. As shown in Table 1, increasing model size consistently improves performance, mirroring scaling trends observed in language and vision domains [brown2020language, dosovitskiy2021an]. Unlike message-passing models that often plateau with scale, G2PM continues to benefit from increased parameterization.
Model Layer. Table 5 shows that deeper encoders enhance performance, while increasing decoder capacity provides limited or negative returns, particularly under the MAE-style masked modeling framework. We attribute this to the simplicity of substructure reconstruction, where expressive decoders can lead to overfitting. Replacing the MAE-style design [he2022masked] with a SimMIM-style variant [xie2022simmim] further degrades performance. This suggests that MAE-style sparsity encourages the encoder to learn stronger context-aware representations, which better align with the inductive bias of graph data.
@@ 3.2 Reconstruction Target
Loss. We also evaluate the impact of loss functions. Replacing $\ell_{2}$ loss with $\ell_{1}$ leads to a notable drop in graph classification accuracy, suggesting that outlier-sensitive objectives better capture nuanced substructure semantics. Additionally, we incorporate anonymous walks [ivanov2018anonymous] following wang2025neural, using a parallel head to reconstruct topological patterns. This auxiliary objective consistently boosts performance, reinforcing the utility of substructures as fundamental modeling units.